OpenStudy (anonymous):

correct me if im wrong What are the x-intercept(s) of the graph of y - x2 = -11x + 24? Answer (-3, 0) and (8, 0) (3, 0) and (-8, 0) I think its this one (-3, 0) and (-8, 0) (3, 0) and (8, 0)

jhonyy9 (jhonyy9):

y=x^2 -11x +24=(x-3)(x-8)

jhonyy9 (jhonyy9):

ok ? can you solve it sure

OpenStudy (anonymous):

it would be c then.

jhonyy9 (jhonyy9):

x_1 =3 x_2 =8

jhonyy9 (jhonyy9):

good luck bye